### v1

# Supporting Bipolar Fabry–Pérot charge interferometer in periodically electron-irradiated graphene

Abstract
Electron optics deals with the wave-nature of charge carriers to induce, investigate and exploit coherent phenomena in solid state devices, in analogy with optics and photonics. Typically, these goals are achieved in complex electronic devices taking advantage of the macroscopically coherent charge transport in two dimensional electron gases and superconductors. Here, we demonstrate a simple counterintuitive architecture employing intentionally-created lattice defects to induce collective coherent effects in the charge transport of graphene. More specifically, multiple Fabry–Pérot cavities are produced by irradiating graphene via low-energy electron-beam to form periodically alternated defective and pristine nano-stripes. The enhanced hole-doping in the defective stripes creates potential barriers behaving as partially reflecting mirrors and resonantly confining the carrier-waves within the pristine areas. The interference effects are both theoretically and experimentally investigated and manifest as sheet resistance oscillations up to 30 K for both polarities of charge carriers, contrarily to traditional electrostatically-created Fabry-Pérot interferometers. Our findings propose defective graphene as an original platform for the realization of innovative coherent electronic devices with applications in nano and quantum technologies.
Keywords: Electron-optics, quantum interference, defective graphene
1 Introduction
Electron optics exploits the wave nature of the electrons in conductors to explore fundamental properties of charges in crystals. Innovative solid state devices and electronic circuits can be developed, in analogy to optical devices and quantum optics experiments, for applications in single-electron quantum optics, nanoelectronics and quantum technologies [1, 2, 3]. Inspired by the optical counterpart, electron optical elements, such as beam splitters [1, 4], focusing lenses [5, 6] and waveguides [1, 7], have been developed to confine and manoeuvre the electron waves. Quantum interference is one of the prototypical manifestations of the wave-like nature of electrons. This phenomenon occurs when two or more electron waves superimpose while coherently travelling in the same conductor. Usually, interference effects are generated in devices where the electrons are spatially confined along specific directions, thus enabling the formation of optics-like interferometers having, for instance, Fabry-Pérot [8, 9, 10] or Mach-Zehnder [11, 12] geometry. Indeed, traditional electronic interferometers have been realized in low dimensional systems [11, 13, 14, 15, 8] or superconductors [16, 17].
The advent of graphene has opened new paths towards novel electron wave devices [18, 19, 20, 21]. Compared to conventional two-dimensional electron gases, graphene offers easily tunable electrical properties, coherent transport accessible up to high temperatures [22], as well as negative refraction of charge carriers [23]. Several optics-like electronic elements, such as interferometers [24, 25], beam splitters [4] and electron-lenses [26, 27], have been currently realized. The key element in graphene coherent devices is the external control of the local electrochemical potential, that allows the formation of energy barriers at the interface of regions of different doping (junction). In the analogy with optics, the electrochemical potential plays the same role of the refractive index in ray optics when considering the semiclassical description [23], whereas the energy barriers act as semitransparent mirrors. Indeed, graphene potential barriers have electron transmission probability () that depends on the junction sharpness and the incident angle (Snell’s law) [28, 29]. In particular, graphene shows Klein tunneling, that is = 1 for orthogonal incidence [20, 29]. Based on these properties, Fabry-Pérot-like resonant effects have been predicted [29] and observed in single or multiple gate graphene field-effect transistors. In particular, these phenomena manifest in the form of gate-induced oscillations in the total charge current flowing along the structure [25, 19, 30, 31, 32, 33, 24, 34]. Indeed, when sweeping the gate voltage (i.e., the electrochemical potential), the Fermi wavelength of the charge carriers is tuned. When the Fermi wavelength matches the wavelengths of the allowed cavity modes, the charge carriers are spatially confined, unlocking resonant phenomena.
Resonant cavities can be realized by large potential steps, i.e., by junctions formed by regions with carriers of opposite polarity ( junctions) or by high/low density of carriers of the same sign ( or junctions). Generally, potential barriers in graphene are created at the metal/graphene interface due to the workfunction difference [19] or by electrostatic gating [35, 20, 36, 37]. The former does not have tunable properties, the latter is effective mainly for junctions. Indeed, Fabry-Pérot (FP) resonances have been predominantly observed in graphene transistors operating in unipolar electronic transport regime [25, 19, 31, 32, 33, 24], that is when the interference occurs only for one sign of the charge carriers.
Moreover, gate-induced potential barriers require complex combinations of different nano-fabrication techniques to create ultra-narrow metal stripes and suffer from fringing electric field effects. The latter cause the enlargement of the effective barrier width and the smoothing of potential step sharpness [30, 37], limiting the lateral resolution of the potential barriers and affecting the interference pattern, especially in multiple gate configuration.
Here, we demonstrate a simple charge carrier interferometer architecture based on the succession of pristine and defective graphene nano-stripes (Figure 1a). In particular, our approach makes use of artificially-created defects in the graphene lattice via electron-beam irradiation to generate charge carrier quantum interference with clear fingerprints in the device charge carrier transport. In other words, typically-undesired defects are here exploited to induce and regulate coherent phenomena in the charge carrier transport of graphene sheets, thus refuting the parallelism between defects and disorder. Moreover, our approach is based on direct writing. It exploits the standard electron-beam lithography system, but it does not involve polymeric resist materials, thus reaching a cavity length scale that is not achievable via standard electrostatic-gating-based approaches.
In our structure, the defects form the mirrors of a FP charge carrier cavity by introducing controlled potential barriers (Figure 1b). Indeed, air exposed defective graphene can be strongly hole doped and operate as potential barrier. Our FP interferometer is realized by regularly alternating parallel defective and quasi-ballistic nano-stripes with nanometric lateral resolution [8]. When tuning the electrochemical potential, periodic oscillations in the sheet resistance of graphene are observed both for holes and electrons, demonstrating the bipolar operation regime. The FP modes are affected by possible disorder in the potential barrier heights, as confirmed by our simulations based on the Landauer-Büttiker formalism. Indeed, the simulations reproduce both the oscillatory behavior of the sheet resistance and the formation of multi-peak resonant modes in the presence of sizeable barrier inhomogeneity. Finally, temperature-dependent measurements show that the FP modes are robust up to 30 K, thus allowing us to extract the charge carrier coherence time, fundamental for the interferometer operation.
2 Results and Discussion
2.1 Fabry-Pérot charge resonances in defective graphene
The FP interferometers are realized in monolayer graphene, which is mechanically exfoliated onto a doped-silicon substrate covered with 300 nm of thermally-grown silicon dioxide. Then, chromium/gold electrical contacts are fabricated by electron-beam lithography followed by thermal evaporation. Afterward, the graphene flakes are shaped in rectangles via reactive ion etching. Finally, the FP interferometer is fabricated by scanning a multi-line pattern with pitch and number of lines onto the graphene sheet with 20 keV electron-beam (see Experimental section for the details and Figure 5, Supporting Information ). An effective linear density of 2 cm-1 defects of different nature is formed in each irradiated line, as estimated by micro-Raman spectroscopy (Discussion 1, Supporting Information) [8]. In general, these defects boost the physical adsorption and activate chemisorption of graphene [39, 40, 41, 42, 8], so that they locally alter the carrier density with respect to pristine graphene. For example, the simple exposure to ambient atmosphere makes defective graphene largely hole-doped [43]. By taking advantage of this simple doping mechanism, we form sharp potential barriers (, where is the Fermi wavevector and is the barrier edge smearing [44, 45]) at the interface between defective and pristine nano-stripes. By keeping the electron-exposed graphene for three days in ambient conditions, the irradiated lines gain large hole-doping of about 300 meV (), whereas the unirradiated areas conserve the as-exfoliated doping of 200 meV (), as extracted via micro-Raman spectroscopy ( Discussion 1, Supporting Information). The barrier height is quantified as meV. By this method, multiple parallel potential barriers are realized with a well-defined pitch of 50 nm and 75 nm [8]. These barriers have Gaussian profile and full-width (, where quantifies also the barrier edge smearing ) of about 25 nm, as imposed by the electron-beam. Hence, the pristine nano-stripes have length 25 nm (for = 50 nm) and 50 nm (for = 75 nm).
The charge transport properties of the FP interferometers are investigated in a four-probe configuration by injecting a constant current () through the device while recording the voltage () for different values of the back-gate voltage () (Figure 1a). We note that this configuration is an accurate counterpart of the optical interferometry experiments, since a constant current keenly simulates the steady intensity of a laser beam. Figure 2a shows the bipolar and periodic oscillations in the resistance vs characteristics measured for a multi-barrier interferometer with = 75 nm and = 133. In particular, the device exhibits several dips superimposed onto an almost standard curve (slightly electron/hole asymmetrical) for graphene.
When calculating the energy-to-resistance transfer function (, where is the energy, i.e. the electrochemical potential), the oscillating behaviour becomes more explicit (Figure 2b). In electron-dominated transport, shows slow oscillations with energy separation of about 13 meV in the energy range close to the Dirac point, whereas it has fast oscillations with 5 meV for 70 meV. This shorter suggests a modification of the FP cavity possibly due to the detailed shape of the potential barrier with energy. Analogous oscillations with similar periodicity occur in the hole transport branch, but, differently from the electron-side, they disappear at -110 meV, i.e. when the electrochemical potential matches the barrier height.
The observed resistance oscillations represent the interference pattern of the electron (hole) waves propagating through the alternating defective and pristine nano-stripes. Indeed, two defective stripes and a pristine one build up a FP cavity (Figure 1b). The 25 nm-long defective stripes act as partially reflecting mirrors capable to transmit phase-coherent carriers [10]. Instead, the pristine zone supports a quasi-ballistic transport, thus it is the medium supporting the interfering carrier-waves (see Discussion 2, Supporting Information). Experimentally, we observe that the periodicity of the resistance dips is independent of the number of barriers (Figure 2c and Figure 9, Supporting Information, for the full transport characterization), indicating that the FP cavities are independent but stringed together. Consequently, the measured interference pattern is averaged over FP interference patterns.
To get more insight into the resonant properties of the graphene FP interferometer, we numerically calculated the sheet conductance when varying the electrochemical potential by combining the envelope-function-based solutions of the Dirac-Weyl equation with the Landauer-Büttiker formalism [47] (see Methods). These simulations assume identical equally-spaced Gaussian barriers with different height and fixed width 2. As shown in Figure 2d, the conductance exhibits several peaks both in the hole- and electron-branch. Moreover, its dependence on energy is stronger for negative energies (hole-side), in full agreement with the experimental observations. In the Landauer-Büttiker approach, peaks in the conductance correspond to the maximum carrier transmission probability of the energy barriers. In our system, this is caused by the constructive interference among the charge waves. Thus, the values of at the conductance peaks (and thus resistance dips) correspond to the energies of the FP modes of the simulated interferometer. Generally, in monolayer graphene the separation energy among the FP modes is defined by the cavity length as , where is the reduced Planck constant and is the Fermi velocity. It is worth noting that in graphene Klein tunneling also affects the FP interference and its visibility: only the waves with non-normal incidence to the barriers (, Figure 1b), but with such that both the transmission and reflection coefficients are reasonably high [25], participate in the resonances. For example, in sharp-edge barriers, 0.5 occurs only when for [44]. The presence of Klein tunneling implies that (i) a quasi-conventional versus characteristics is superimposed to the FP oscillations, as shown in Figure 2a; (ii) the effective is longer than the distance between the potential barriers, i.e., along the orthogonal direction.
The equally spaced FP modes of our simulations show a good qualitative and quantitative agreement with the experimental data. In particular, the separation energy of the simulated modes is 25 meV (the same order of the experimental 13 meV), corresponding to 85 nm and an effective incidence angle that is (see Discussion 3, Supporting Information ). By exploiting the same argument, 160 nm and . It is worth noting that the simulations consider identical barriers and do not take into account the possible extension of the defects into the pristine stripes due to lattice re-arrangement, as it may occur in the experiments. These defects contribute to the quasi-ballistic transport as elastic (phase-preserving) scattering centers, making the cavity length longer compared to a defect-free stripe. Consequently, the longer the , the smaller the . Moreover, due to elastic scattering, the incidence angle at the barriers step can be also modified, and thus the barriers transmittance, influencing the interference pattern visibility.
2.2 Effects of potential disorder and temperature
When reducing the pitch down to 50 nm (thus 25 nm), distinct oscillations in the sheet resistance are still clearly observable (see Figure 3a), thus confirming the robustness of the proposed method for obtaining FP interferometers. In this case, the numerical simulations predict FP modes every 40 meV, corresponding to 55 nm and an effective (similar to the cavity length/normal barrier distance ratio and of = 75 nm).
The analysis of reveals that in the electron-branch the slow oscillations with 20 meV (violet arrows in Figure 3b, corresponding to 105 nm and ) are split into fast oscillations with 4 meV (green arrows in Figure 3b). Instead, in the hole-branch, only the fast oscillations can be clearly distinguished. This behaviour is reproduced also when probing a different number of barriers (see Figure 11b, Supporting Information).
In order to understand the origin of these fast oscillations, we simulated the sheet conductance including randomly-distributed disorder in the potential barrier height (Figure 12, Supporting Information). Indeed, experimentally potential inhomogeneities are unavoidable. This disorder can originate from the substrate electrostatics, carrier puddles, or local doping inhomogeneities within the defective stripes [48]. The latter are more probable at short pitches because of the higher interaction among the irradiated stripes [8]. The simulations consider different degrees of potential disorder (see Figure 3c). As shown in Figure 3d, this disorder strongly impacts the FP resonances. Differently from the ideal case (where the resonances are a set of well-defined spectral bands), the FP modes are less visible, when slowly increasing the potential fluctuations (disorder 1 - 2). For higher potential disorder (disorder 3 - 4), the main FP resonances are broader and less resolved. Moreover, they show superimposed fast oscillations with 5 meV (inset in Figure 3d). Tentatively, these multi-peak resonances are explained by the fact that the disorder locally induces a slight variation of the potential barrier height . Consequently, the FP modes energy is also modified due to their spectral dependence on , as visible by comparing the curves at slightly different in Figure 2d. Hence, the main multi-peak resonances arise from the spectral overlap of several FP modes having close resonance energies. For highly disordered barriers, instead, the main resonant modes and the fast oscillations almost completely disappear (disorder 5).
Our graphene FP interferometers show resonant oscillations of resistance up to a temperature = 30 K (see Figure 4). At all , the resistance oscillations manifest at the same energies. By rising , they are slowly suppressed, as expected by the thermal reduction of the charge carrier coherence length (), until they disappear when becomes comparable to . The coherence length is defined as , with the diffusion constant, the Fermi velocity, the mean-free-path and the coherence time. In our interferometers, is about 0.02 m2/s, since the lower bound of the mean-free-path is 30 nm (see Discussion 2, Supporting Information). Consequently, imposing ( = 30 K) , the maximum value of at 30 K can be quantified in 0.7 ps for = 50 nm and 2 ps for = 75 nm. It is worth noting that the longer for = 75 nm is reflected in the fact that the oscillations are not fully suppressed at 30 K, thus confirming our analysis.
Finally, we stress that the FP interference pattern is reproducible over different thermal cycles, even the multi-peak resonance modes for = 50 nm (see Figure 11, Supporting Information). This feature fully confirms the quantum interference origin of the observed resistance oscillations and rules out other trivial explanations.
3 Conclusions
We presented an original approach to generate coherent quantum phenomena in graphene sheets by exploiting the wave nature of charge carriers. In particular, electron-irradiation-induced defects in the graphene lattice are exploited to mediate the coherent carrier transport by acting as partially transmitting mirrors, that confine the charge carriers within pristine quasi-ballistic nano-stripes (resonant structures). The regular alternation of defective and pristine nano-stripes creates independent FP cavities, whose interference fringes manifest as thermally-robust (up to 30 K) oscillations in the graphene sheet resistance. These oscillations are observed in series of FP interferometers having different cavity lengths. Moreover, the charge carrier interference patterns are present for both electrons and holes, revealing a bipolar behaviour, that is typically hardly observable in gate-induced barriers. As confirmed by numerical simulations, possible disorder in the barrier height generate multi-peak main FP modes. Finally, the interference pattern is not correlated to the number of the potential barriers, indicating that the FP cavities are independent.
Differently from the standard electrostatic approach based on gating, this methodology is extremely simple. On the one hand, it avoids the use of metal stripes that limit the minimum cavity length (typically 50 nm), offering the possibility to fully scale down and produce resonant structures also on a chip scale-level (if using high quality large-area graphene). On the other hand, thanks to the enhanced chemical reactivity, it offers an additional knob to control the resonant cavities: when properly functionalized [40], the defective graphene allows to define ad hoc the properties of the potential barriers. Consequently, electron-irradiated graphene represents an original route to develop innovative coherent electronics devices for the manipulation of carrier waves and the manifestation of quantum charge-optics effects.
4 Experimental Section
Graphene interferometers fabrication
Monolayer graphene flakes are mechanically exfoliated (natural graphite from NGS Trade and Consulting GmbH and blue tape from Nitto Italia srl) onto doped-silicon substrates covered with 300 nm of thermally-grown silicon dioxide. The substrate chips are initially cleaned by oxygen plasma at 100 W for 5 minutes to promote the graphene adhesion and remove organic residues from the substrate surface to maximize the defect patterning lateral resolution [4].
First, electrical contacts to the graphene sheet are fabricated by combining electron-beam lithography with thermal evaporation of a 5 nm/50 nm-thick film of chromium/gold. The graphene flakes are subsequently dry-etched via reactive ion etching method with a polymer mask to obtain the rectangular shape (10 m-wide and different length from 2 m to 10 m). Finally, the chips are deeply cleaned by resist remover (All resist, AR 600-71) to eliminate the residues from the graphene surface.
The defective nano-stripes are produced by scanning a multi-line pattern on the graphene surface with an electron-beam at 20 keV. The pitch between the defective lines is = 50 nm and = 75 nm, whereas the irradiation step-size along each line is 7.8 nm. The number of defective stripes are = 40, 200 for = 50 nm and = 67, 107, 133 for = 75 nm. The electron beam is produced by a commercial lithography system (from Raith GmbH) and it has a beam spot of 25 nm. For all patterns, the electron-beam current is about 100 pA, delivering a linear dose of 31 mC/cm and resulting in a dwell-time of 250 s.
Prior to the electrical measurements, the irradiated samples are exposed to ambient atmosphere for three days in order to hole-dope the defective stripes and establish the potential barriers.
The crystal lattice of the irradiated-graphene is studied in ambient by micro-Raman spectroscopy (Renishaw). The Raman spectrum is acquired by scanning the graphene surface with a laser at 532 nm with a 100 objective (NA = 0.85), that corresponds to a lateral resolution 1 m. The map step-size is 500 nm. The laser power is set at 118 W to exclude any possible laser heating of the lattice. More details about the defective multi-stripe topography and crystallographic structure are reported in Ref. [8].
Charge transport measurements
The interferometric properties of the periodically-irradiated graphene is measured in a dry cryostat (OptiStat Dry by Oxford Instruments), having a base temperature of 2.5 K. The measurements as a function of temperature are carried out by employing a temperature controller (Mercury iTC by Oxford Instruments), which allows the sweep from 2.5 K up to 300 K. The temperature stability of the system is better than 50 mK.
The electronic transport of the carrier-interferometers is measured in a four-probe configuration. An AC source-to-drain current at 17 Hz is injected into the graphene and the voltage drop across two probe contacts is measured via lock-in technique (Stanford SR830). The DC gate voltage (sweep step of 100 mV) is applied by a source-meter (Keithley 2614B) in back-gate configuration.
Charge transport simulations
In order to keep the computational time within reasonable limits but without loss of generality, we simulated a 1 m-wide graphene sheet containing 10 tunnel barriers, that are orthogonal to the transport direction. The barriers have a Gaussian profile along the transport direction with peak amplitude in the range 80-150 meV (10 meV step) and full-width 2 25 nm. We assume the distance between the barrier peaks to be either 50 nm or 75 nm.
The simulations are performed by combining the solution of the Dirac-Weyl envelope-function equation with the Landauer-Büttiker formalism [35, 50, 47].
In the low-energy range, the graphene electron-wave function can be written in terms of four envelope functions, , , , , each one corresponding to one of the two triangular sublattices and and of the two unequivalent Dirac points and . These four envelope functions have to satisfy the Dirac-Weyl equation, with Dirichlet boundary conditions enforced at the sheet edges.
The transport problem is turned into a set of much simpler ones using a recursive scattering matrix approach. The graphene sheet is partitioned into a series of slices, that are as wide as the graphene sheet (transverse direction, ), but with the length along the transport direction () sufficiently short to make the variation of the potential energy along the -direction negligible within each slice. As a consequence, within each slice the longitudinal wave vector is conserved and each envelope function can be factored into a part , that depends only on the transverse coordinate , and into a plane wave in the transport direction. Therefore, within each slice the Dirac-Weyl equation becomes the following differential eigenproblem:
where , , and are the Pauli matrices, , (with the graphene Fermi velocity, the reduced Planck constant, the electron energy), , and is the effective width of the sheet. In order to optimize the calculations and avoid the fermion doubling problem, which arises from a standard direct space discretization of the Dirac equation, the problem is transformed into this equivalent one, as follows
where the function is defined in the range as
The new differential problem (Equation 2) has periodic boundary conditions on and thus it can be numerically solved (with a proper frequency cut) in the reciprocal space, obtaining a basis of solutions of the Dirac-Weyl equation. These solutions represent the transport modes in the slice.
Once the transmission modes of two adjacent slices are computed, the scattering matrix connecting these modes is obtained by mode-matching technique. For each mode, we write the total wave function (in terms of the unknown reflection and transmission coefficients), that results from the injection of that single mode into the pair of slices. This wave function has to satisfy the continuity condition at the interface between the two slices. By projecting the resulting set of continuity equations onto a basis of sine functions, a linear system in the reflection and transmission coefficients is thus obtained. Solving this system, the scattering matrix of the two coupled slices is calculated. Recursively composing the scattering matrices that connect all the slices in the sheet, the overall scattering matrix, and, in particular, the transmission matrix between the modes in the input and output leads, is finally computed. From this matrix, the conductance of the graphene can be then estimated by exploiting the Landauer-Büttiker formula as
where is the elementary charge and is the Plank constant. The sum is performed over all the modes that propagate in the input and in the output leads, which are specified by the indices and , respectively.
Finally, the disorder in the potential height is reproduced [50] by adding to the potential profile a sum of two-dimensional Gaussian functions, i.e., , where is the distance from the Gaussian center and = 35.7 nm. These Gaussian functions are randomly distributed over the graphene sheet with a concentration equal to cm-2 and differ in peak amplitude (). is randomly chosen, with a uniform distribution, between and meV.
Supporting Information
Supporting Information is available from the Wiley Online Library or from the author.
Acknowledgements
The authors thank S. Cassandra for fruitful discussions. The authors wish to acknowledge C. Coletti from the Istituto Italiano di Tecnologia for the access to the micro-Raman facility.
F. B. and F. P. acknowledge partial financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call PRIN 2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU – EQUATE Project, "Defect engineered graphene for electro-thermal quantum technology" - Grant Assignment Decree No. 2022Z7RHRS.
F. B. acknowledges partial financial support by the MUR - Italian Minister of University and Research under the “Research projects of relevant national interest - PRIN 2020” - Project No. 2020JLZ52N, title “Light-matter interactions and the collective behavior of quantum 2D materials (q-LIMA)”.
S.R. acknowledges partial financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call PRIN 2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU – TRUST project "Trampolines as ultra-Sensitive thermomechanical bolometers" - Grant Assignment Decree No. 2022M5RSK5.
P. M. and M. M. acknowledge partial financial support by the Italian Ministry of University and Research (MUR) in the framework of the FoReLab and CrossLab project (Department of Excellence) and by the European Union, Next-GenerationEU, National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment N. 1.4, CUP N. I53C22000690001, through the National Centre for HPC, Big Data and Quantum Computing (“Spoke 6: Multiscale Modelling Engineering Applications” and “Spoke 10: Quantum Computing”).
Conflict of Interest
The authors declare no conflict of interest.
Supporting information:
Bipolar Fabry–Pérot charge interferometer in periodically electron-irradiated graphene
Figure S1: Description of the graphene periodical irradiation
Discussion 1: Doping and defect probing via micro-Raman spectroscopy
A quantitative analysis of the charge doping and the density of defects induced in the lattice via electron-irradiation can be extrapolated from micro-Raman spectroscopy data.
Doping probing
In the low defect density regime, the correlation plot method from Ref. [1] is valid [2, 3] and thus, the charge doping can be estimated. As reported in Figure 6, the two-dimensional (2D) carrier density is quantified as cm-2 for pristine graphene (from the as-exfoliated values [4, 5]), cm-2 for = 75 nm, and cm-2 for = 50 nm.
Defects probing
As shown in Figure7a-b, the Raman spectrum exhibits the and bands. These bands are the unambiguous signature of defects in the crystal and their analysis gives insight into both the nature [6] and the density of the defects [7].
The nature of the defects in the crystal can be identified from the ratio between the and the peak intensities (). The / 11 (for = 75 nm) and 9 (for = 50 nm) indicates mostly sp3-like defects, as expected for defective graphene exposed to air atmosphere for several days [5, 6].
Instead, the two-dimensional (2D) sheet density of defects () is computed correlating the intensities of the D and G peaks [ and , respectively]. According to the formula [7]
where is the pump laser energy (in our case, 2.33 eV), we obtain a 2D defect density of for nm [uniformly distributed 1.05, Figure 7c] and for nm [uniformly distributed 1.45, Figure 7d]. As discussed in the main text, the samples consist of periodic repetitions of defective and pristine nano-stripes. The laser spot size is 1 m, so that it cannot spatially resolve the defective and pristine areas. This means that the acquired Raman spectrum is an average spectrum containing information from multiple defective and pristine nano-stripes. To give an estimation of the effective density of defects along each nano-stripe, we consider that, for a given pitch, the number of lines per unit length is . Therefore, can be defined in terms of one-dimensional (1D) density as follows
Consequently, combining Eq. 5 with Eq. 6, we estimate an effective 1D density of defects and for nm and 50 nm, respectively.
Interestingly, the values of are not constant [although all the lines are exposed with the same beam parameters (see Methods in the main text)], indicating that the dependence on the pitch is different from a simple . This is explained by the presence of a cross-talk among the defective nano-stripes. This cross-talk arises from proximity effects occurring during the irradiation process, as explained more in details in Ref. [8], where the phenomenological model describing the dependence of the values on the pitch is introduced, that is
Figure S4: FP resonances at different source-drain currents
Discussion 2: Phase-coherent quasi-ballistic transport through periodically irradiated graphene sheet
For = 75 nm (50 nm), the pristine stripes have a length of 50 nm (25 nm). Here, a coherent quasi-ballistic transport can be safely assumed. Generally, the mean-free path () can be estimated from the mobility, which is extracted by fitting the total measured sheet resistance [9]. When focusing on the more symmetric hole-dominated transport (Figure 2.a and 3.a in the main text), the mean-free path is quantified as high as 30 nm. This is a lower bound value, as the measured resistance arises from the series of the defective and pristine graphene resistances. Consequently, quasi-ballistic trajectories are preserved, ensuring phase-matching among the carriers that travel back and forward within the pristine stripes after the reflection at the potential step built at the interface with the defective stripes.
In the defective stripes, quantum interference effects are excluded. These stripes only take part in creating the mirrors, while quasi-ballistically transmitting phase-coherent carriers into the next nearest FP cavity. Based on the Raman spectroscopy results, the disorder in defective stripes is in the low density regime [2, 3]. In this regime, the carrier phase coherence length is almost unaffected, as demonstrated in similar electron-irradiated graphene [10]. Instead, the mean-free-path becomes shorter compared to the pristine lattice [10], confirming our underestimation of in the pristine stripes.
Figure S5: Carrier transport through different number of potential barriers
Discussion 3: Charge carrier effective incident angle at the potential step
As explained in the main text, the presence of Klein tunneling imposes that the FP resonances are visible only when charge carriers have an incident angle at the potential step such that the transmission and reflection coefficients are maximized. In order to give the estimation of an effective both in the experiments and in the simulations, we assume a simplified picture (Figure 10) and we apply some trigonometrical calculations. Considering the ideal case where the pristine zone has length equal to = - 2 and by using the cavity length extracted from , the incident angle is quantified as
Figure S7: Thermal cycle effects on the FP resonances
Figure S8: Randomly-distributed disorder in the potential barrier height